OpenStudy (anonymous):

A dolphin emits ultrasound at 100kHz and uses the timing of reflections to determine the position of objects in the water. What is the wavelength of this ultrasound? Assume that temperature of water is 20C.

OpenStudy (anonymous):

velocity = frequency * wavelength Use velocity of sound in water at the given temperature

OpenStudy (shamim):

u know the velocity of sound is v T= temperature in kelvin scale then the relationship between v and T is

OpenStudy (shamim):

\[v \alpha \sqrt{T}\]

OpenStudy (shamim):

so\[\frac{ v1 }{ v2 }=\sqrt{\frac{ T1 }{ T2 }}\]

OpenStudy (shamim):

here \[T1=293K, T2=273K, v2=332m/\sec ^{2}\]

OpenStudy (shamim):

r u following me